A 90 m long train is sitting at the beginning of the tracks in a railway yard. The train begins to a celebrate uniformly from its start position. The very front of the engine has a velocity of 20 m/s when it passes of railway worker who is standing at the first switch 180 m away from the very front of the engine when it started moving. calculate the velocity of the end of the very last car as it passes the workers standing on the tracks.


1
Expert's answer
2021-03-04T15:18:28-0500
"v^2=2as\\20^2=2(180)a\\a=1.1\frac{m}{s}\\u^2-v^2=2ad\\u^2-20^2=2(1.1)(90)\\u=24.5\frac{m}{s}"